alexandra works for two clients during each month. she works a total of 190 hours. during the month. she works 63 hours more with client y than for client x. how many hours does she work for each client?
1
Expert's answer
2020-05-01T18:23:25-0400

x + y = 190

y = x + 63


x + x + 63 = 190

2x = 127

x = 63.5

y = 126.5

Answer: x = 63.5 hours; y = 126.5 hours.